OpenStudy (anonymous):

Solve 3x^2 + 4x = –5 Show your work please.

jhonyy9 (jhonyy9):

D=16-60=-44 x_1,2=(-4+/- sqrt(-44))/12 = (-4+/- isqrt(4*11))/12 = =(-4+/- 2isqrt11)/12 =(-2+/- isqrt11)/6

OpenStudy (anonymous):

u bring the -5 to the left side,u get 3x^2+4x+5, then u solve for (delta) delta= b^2-4.5 u get -44 and since delta<0 there are no solutions

OpenStudy (anonymous):

sorry delta= b^2-4.5.3

jhonyy9 (jhonyy9):

yes you know it right but if delta is less than zero the function will have complex roots

jhonyy9 (jhonyy9):

- so how i have got there

jhonyy9 (jhonyy9):

ok ?